The speed of a moving bullet can be determined by allowing the bullet to pass through
two rotating paper disks mounted a distance
69 cm apart on the same axle. From the
angular displacement 43.5◦
of the two bullet holes in the disks and the rotational speed
368 rev/min of the disks, we can determine
the speed of the bullet.
What is the speed of the bullet?
Answer in units of m/s.
Don't round answer
1
Expert's answer
2019-12-19T11:37:27-0500
"\phi=43.5\frac{\pi}{180}=0.76\ rad"

"n=\frac{368}{60}=6.13\frac{rev}{s}"

"t=\frac{0.76}{2\pi(6.13)}=0.02\ s"

"v=\frac{0.69}{0.02}=34.5\frac{m}{s}"